### v2

# Accelerated Sinkhorn Algorithms for Partial Optimal Transport

Abstract
Partial Optimal Transport (POT) addresses the problem of transporting only a fraction of the total mass between two distributions, making it suitable when marginals have unequal size or contain outliers. While Sinkhorn-based methods are widely used, their complexity bounds for POT remain suboptimal and can limit scalability. We introduce Accelerated Sinkhorn for POT (ASPOT), which integrates alternating minimization with Nesterov-style acceleration in the POT setting, yielding a complexity of $\mathcal{O}(n^{7/3}\varepsilon^{-5/3})$. We also show that an informed choice of the entropic parameter $\gamma$ improves rates for the classical Sinkhorn method. Experiments on real-world applications validate our theories and demonstrate the favorable performance of our proposed method. Our code is open-source at the link github.
Index Terms—  Partial Optimal Transport, Gradient-based optimization, Nesterov’s momentum.
1 Introduction
Optimal Transport (OT) has been a core tool to compare probability distributions, with broad impact from economics to machine learning [11, 12, 18, 23]. Its success stems from efficient computational methods [3, 10, 11, 20] and the role of entropic regularization, which improves statistical properties and enables seamless integration into deep learning frameworks [2, 10, 12, 21, 5, 7]. Classical OT, however, requires equal mass and is sensitive to outliers [29]. POT addresses these issues by relaxing mass constraints, making it more robust in practical tasks such as image analysis, few-shot learning and distributional learning [35, 36, 34, 17, 22, 6]. However, existing computational approaches remain unsatisfactory. A practical path to POT is to reduce it to a balanced OT instance by adding dummy nodes and then apply Sinkhorn‐type iterations [6]. While straightforward, naive variants can be infeasible [17]. A feasible formulation with guarantees was established in [22], yielding a Sinkhorn complexity of $\mathcal{O}\left(n^{2}/\varepsilon^{4}\right)$ and proposing APDAGD with $\mathcal{O}\left(n^{2.5}/\varepsilon\right)$. However, applied researchers note that tuning the Sinkhorn parameter often yields stronger performance, while APDAGD’s extra $\mathcal{O}(n^{0.5})$ factor limits its scalability. Thus, our central questions are:
  1. 1.
    Is there a first-order method for entropic POT with a better $\varepsilon$-dependence than feasible Sinkhorn and a better $n$-dependence than APDAGD?
  2. 2.
    How does varying the regularization parameter $\gamma$ affect the feasible Sinkhorn for POT, and can parameter tuning improve both theory and practice?
We summarize our contributions as follow:
  1. 1.
[…]
  2. 2.
    We analyze the classic Sinkhorn, showing how the careful choice of $\gamma$ improves both theory and practice, achieving a better complexity $\mathcal{O}\left(n^{2}\|C\|^{2}_{\infty}/\varepsilon^{\frac{3p+1}{p}}\right)$. Here $p\in[1,\infty)$ is a tunable analysis parameter. We write $o(1)$ to denote a quantity that can be made arbitrarily small by taking $p$ sufficiently large.
  3. 3.
    Through experiments on color transfer and point cloud registration, we show that ASPOT converges faster and produces higher-quality solutions than existing baselines, and that tuning $\gamma$ further boosts Sinkhorn’s performance.
Algorithm Complexity
Feasible Sinkhorn [22] $\mathcal{O}\!\left(n^{2}\|C\|_{\infty}^{2}/\varepsilon^{4}\right)$
APDAGD [22] $\mathcal{O}\!\left(n^{2.5}\|C\|_{\infty}/\varepsilon\right)$
ASPOT (This paper) $\mathcal{O}\!\left(n^{7/3}\|C\|_{\infty}^{4/3}/\varepsilon^{5/3}\right)$
Tuned Sinkhorn (This paper) $\mathcal{O}\!\left(n^{2}\|C\|_{\infty}^{2}/\varepsilon^{3+o(1)}\right)$
Table 1: Complexity of POT algorithms
2 Preliminaries
2.1 Problem formulation
Notation: Vectors are bold lowercase, matrices bold uppercase, and $\mathbf{1}_{n}$ is the all-ones vector in $\mathbb{R}^{n}$. For any vector or matrix $X$, $X^{\top}$ denotes transpose, $\langle A,B\rangle=\sum_{i,j}A_{ij}B_{ij}$, and $H(x)=-\langle x,\log x\rangle$. We use $\|x\|_{1}=\sum_{i}|x_{i}|$, $\|A\|_{1\to 2}=\max_{\|x\|_{1}=1}\|Ax\|_{2}$, $\rho(a,b)=b-a+a\log(a/b)$ for $a,b>0$, and $\operatorname{diag}(x)$ for the diagonal matrix with $x$ on its diagonal. Dual variables are $z=(u,v,w)$ with $\nabla$ their derivatives. Algorithmic complexity is denoted $\mathcal{O}(\cdot)$.
Firstly, recall the POT problem. Given two discrete distributions $r,c\in\mathbb{R}^{n}_{+}$ with possibly different masses and a transport budget $s\leq\min\{\|r\|_{1},\|c\|_{1}\}$, the goal is to find a transport plan $X\in\mathbb{R}^{n\times n}_{+}$ minimizing the cost
$\text{POT}(r,c,s)=\min_{X\in\mathcal{U}(r,c,s)}\langle C,X\rangle$ (1)
where $C\in\mathbb{R}^{n\times n}_{+}$ is the cost matrix and the feasible set is $\mathcal{U}(r,c,s)=\{X\geq 0:X\mathbf{1}_{n}\leq r,\,X^{\top}\mathbf{1}_{n}\leq c,\,\mathbf{1}_{n}^{\top}X\mathbf{1}_{n}=s\}.$
Following the success of entropic OT, entropic regularization leads to $\text{min}_{X\in\mathcal{U}(r,c,s)}\langle C,X\rangle-\gamma H(X)$, where $\gamma>0$ is the regularization parameter. The corresponding dual [22] introduces potentials $u,v\in\mathbb{R}^{n}$ and $w\in\mathbb{R}$, giving
$\min_{z}\varphi(z)=\|B(z)\|_{1}$ $+\sum_{i=1}^{n}\exp(u_{i})+\sum_{j=1}^{n}\exp(v_{j})$
$-\langle u,r\rangle-\langle v,c\rangle-ws,$ (2)
where $B(z):=mD_{1}KD_{2}$, with $D_{1}=\mathrm{diag}(e^{u_{i}})$, $D_{2}=\mathrm{diag}(e^{v_{j}})$, $K=e^{-C/\gamma}$, and $m=e^{w}$.
Next, we define the standard notion of $\varepsilon$-approximation:
Definition 1. For $\varepsilon>0$, a transport plan $X$ is an $\varepsilon$-approximation of POT($r,c,s$) if $X\in\mathcal{U}(r,c,s),\langle C,X\rangle\leq\langle C,X^{\star}\rangle+\varepsilon,$ where $X^{\star}$ is an optimal transport plan of the POT problem.
3 Algorithm Development
Both proposed methods share a two-stage structure where in the first stage, we design a solver tailored to the entropic-regularized POT problem. In the second stage, we invoke the Approximating POT procedure (Algorithm 2, Lines 1–8) together with the ROUND-POT routine, which ensures an $\varepsilon$-approximate solution that remains feasible. Coupled with our ASPOT and Tuned Sinkhorn variants, this framework achieves improved complexities of $\mathcal{O}(n^{7/3}\varepsilon^{-5/3})$ and $\mathcal{O}(n^{2}/\varepsilon^{(3p+1)/p})$, respectively. A full comparison of POT complexities for existing and proposed methods is summarized in Table 1.
Algorithm 1 Greenkhorn Step(u,v,w)(u,v,w)
1:if $t\bmod 3=0$ then
2:  $u\leftarrow u+\log r-\log\!\big(r(B(u,v,w))+e^{u}\big)$
3:else if $t\bmod 3=1$ then
4:  $v\leftarrow v+\log c-\log\!\big(c(B(u,v,w))+e^{v}\big)$
5:else
6:  $w\leftarrow w+\log s-\log\|B(u,v,w)\|_{1}$
7:end if
8:return $(u,v,w)$
Theorem 1.
Algorithm 2 returns an $\varepsilon$-approximate transport plan in $\tilde{O}\left(\frac{n^{7/3}\|C\|_{\infty}^{4/3}(\log(n))^{1/3}}{\varepsilon^{5/3}}\right)$ arithmetic operations.
Remark 1.
Theorem 3 establishes the iteration complexity required for Algorithm 2 to reach an $\varepsilon$-approximate solution. Since each iteration costs only $\tilde{O}(n^{2})$ arithmetic operations, this yields the overall complexity stated in Theorem 1.
Algorithm 2 Accelerated Sinkhorn POT
1:Marginals $r,c$, cost matrix $C$, transport mass $s$, accuracy $\varepsilon$
2:Set $\gamma=\varepsilon/(4\log n)$, $\tilde{\varepsilon}=\varepsilon/(8\|C\|_{\max})$
3:if $\|r\|_{1}>1$ then
4:  $\tilde{\varepsilon}\leftarrow\min\{\tilde{\varepsilon},\,\tfrac{8(\|r\|_{1}-s)}{\|r\|_{1}-1}\}$
5:end if
6:if $\|c\|_{1}>1$ then
7:  $\tilde{\varepsilon}\leftarrow\min\{\tilde{\varepsilon},\,\tfrac{8(\|c\|_{1}-s)}{\|c\|_{1}-1}\}$
8:end if
[…]
10:Initialize $t=0$, $\theta_{0}=1$, $\tilde{z}^{0}=\check{z}^{0}=\mathbf{0}_{n+1}$.
11:while $E_{t}>\tilde{\varepsilon}$ do
12:  $\bar{z}^{t}\leftarrow(1-\theta_{t})\check{z}^{t}+\theta_{t}z^{t}$
13:  $z^{t+1}\!\leftarrow\!\bar{z}^{t}-\frac{\gamma}{3(\|r\|_{1}+\|c\|_{1}-s)\theta_{t}}\nabla\varphi(\bar{z}^{t})$
14:  $\grave{z}^{t}\leftarrow\bar{z}^{t}+\theta_{t}\,(z^{t+1}-z^{t})$
15:  $\hat{z}^{t}\leftarrow\textsc{GreenkhornStep}(\grave{z}^{t})$
16:  $z^{t}\leftarrow\arg\min\{\varphi(z):z\in\{\check{z}^{t},\hat{z}^{t}\}\}$
17:  $\check{z}^{t+1}\leftarrow\textsc{GreenkhornStep}(z^{t})$
18:  $\theta_{t+1}\leftarrow\frac{\theta_{t}\big(\sqrt{\theta_{t}^{2}+4}-\theta_{t}\big)}{2}$, $t\leftarrow t+1$
19:end while
20:Form primal: $\widetilde{X}\leftarrow B(z^{t})$ with $B_{ij}(z)=\exp\!\big(-C_{ij}/\gamma+u_{i}+v_{j}+w\big)$
21:Round $\widetilde{X}$ to $\bar{X}$ using ROUND-POT[22]
22:Output: $\bar{X}$
3.1 Acceleratd Sinkorn POT Algorithm
Algorithm 2 is designed to combine two ideas: the momentum mechanism of Nesterov and the fast per-iteration progress of Greenkhorn in low-precision regimes. A naive combination, such as the Alternating Accelerated Method (AAM) [1], leads to an additional $\mathcal{O}\left(n^{0.5}\right)$ multiplicative factor in the complexity, which makes it impractical for large-scale problems in practice. Instead, we let the Greenkhorn update serve directly as the coupling step within Nesterov’s framework in line 12. After each extrapolation, the algorithm applies the most effective coordinate correction, stabilizing the momentum and ensuring that progress concentrates on the largest residual (line 14). To monitor this, we use a function measuring the error per iteration:
$E(t)=\big|\mathbf{1}_{n}^{\top}B^{t}\mathbf{1}_{n}-s\big|+\|B^{t}\mathbf{1}_{n}+\tilde{p}-r\|_{1}+\|(B^{t})^{\top}\mathbf{1}_{n}+\tilde{q}-c\|_{1},$
Equally important is the descent of the dual objective 2.1. The Greenkhorn updates in lines 14 and 16 enforce $\varphi(\grave{z}^{t})\!\geq\!\varphi(\hat{z}^{t})$ and $\varphi(z^{t})\!\geq\!\varphi(\check{z}^{t+1})$, ensuring steady decrease for 2.1.
Fig. 1: Color Transfer Experiment
3.2 Tuned Sinkhorn Technique
Motivated by recent insights in the OT literature [14], we revisit the role of the regularization parameter $\gamma$ in the Sinkhorn algorithm for POT [22]. By scaling $\gamma$ relative to the minimal entropy term $H_{\min}(r,c)=\min\{H(r),H(c)\}$ and the accuracy parameter $\varepsilon$, the classical Sinkhorn method attains an improved complexity bound of $\mathcal{O}\left(n^{2}H_{\text{min}}(r,c)^{1/p}\|C\|^{2}_{\infty}/\varepsilon^{\frac{3p+1}{p}}\right)$. This refined analysis clarifies the performance of Sinkhorn in practice and extends its theoretical guarantees to the partial transport setting.
Theorem 2.
The feasible Sinkhorn algorithm [22] for POT with the chosen regularization $\gamma=\sqrt[p]{2\varepsilon/49H_{\text{min}}(r,c)}$ and $p\in[1,\infty)$ returns the $\varepsilon$-approximation in
$\mathcal{O}\left(n^{2}H_{\text{min}}(r,c)^{1/p}\|C\|^{2}_{\infty}/\varepsilon^{3+o(1)}\right)$ (3)
arithmetic operations.
Proof sketch.
The error $\langle\bar{X}-X^{\star},C\rangle$ decomposes into rounding, regularization, and the gap between the extended OT solution and the POT optimum. The rounding error is guaranteed by ROUND-POT in Algorithm 2 [22], contributing $O(\varepsilon^{\prime})$. The regularization error is controlled by the entropy-regularized OT analysis [14], which also applies to POT since the extra (n+1,n+1)(n{+}1,n{+}1) entries scale with $1/\gamma$ to ensure stability. Finally, Proposition 1 in [6] bounds the remaining gap within $O(\varepsilon^{\prime})$. Combining yields $\langle\bar{X}-X^{\star},C\rangle\leq\varepsilon+\tilde{O}(\varepsilon^{2})$. Substituting $\varepsilon^{\prime}=H_{\min}(r,c)\gamma^{p}$, $A=\tilde{O}(\|C\|_{\infty}/\varepsilon)$, and $\gamma=\sqrt[p]{2\varepsilon/(49H_{\min}(r,c))}$, together with the $O(n^{2})$ per-iteration cost of Sinkhorn, completes the argument. ∎
3.3 Theoritical lemmas and theorems
We refer the reader to [[30], Appendix] for the construction of the matrix $A$ (via slack variables turning POT into a linear programm), complete proofs, extra algorithms, and further experiments. From [22], we recall the bound on the dual variables $z^{\star}$:
$R:=\frac{\|C\|_{\infty}\max(\|r\|_{1},\|c\|_{1})}{\gamma(\max(\|r\|_{1},\|c\|_{1})-s)}-\log\!\Big(\min_{1\leq i,j\leq n}(r_{i},c_{j})\Big).$ (4)
Lemma 1.
Define the Lipschitz constant $L=\frac{\|A\|^{2}_{1\rightarrow 2}}{\mu_{f}}$ where $\mu_{f}=\frac{\gamma}{\|r\|_{1}+\|c\|_{1}-s}$. Let $\delta_{t}=\varphi(\check{z}^{t})-\varphi(z^{\star})$, where $z^{\star}$ is an optimal dual solution of the entropic-regularized POT problem. Then we have
$\delta_{t+1}\leq(1-\theta_{t})\delta_{t}+\tfrac{L}{2}\theta_{t}^{2}\Big(\|z^{\star}-\tilde{z}^{t}\|^{2}-\|z^{\star}-\tilde{z}^{t+1}\|^{2}\Big).$
Proof sketch.
The key step to derive $\delta_{t}$ is to couple the extrapolated iterate $\bar{z}^{t}$ with both the new accelerated point $\tilde{z}^{t+1}$ and the previous $\tilde{z}^{t}$, which produces a telescoping quadratic term $\|z^{\star}-\tilde{z}^{t}\|^{2}-\|z^{\star}-\tilde{z}^{t+1}\|^{2}$. This coupling makes the quadratic distance terms cancel in a telescoping way, so the error can only get smaller over time. Finally, the Greenkhorn corrections enforce a monotone descent of the dual objective, namely $\varphi(\grave{z}^{t})\!\geq\!\varphi(\hat{z}^{t})\!\geq\!\varphi(z^{t})\!\geq\!\varphi(\check{z}^{t+1})$, allowing the recurrence to be transferred to the actual iterates. ∎
Remark 2.
Lemma 1 allows us to express $\delta_{t+1}$ recursively in terms of $\delta_{t}$ and a decreasing potential involving $\|z^{\star}\|$. Next, from the definition of $\theta_{t}$ implies the relation $\tfrac{\theta_{t+1}}{\theta_{t}}=\sqrt{1-\theta_{t+1}}$. Hence, by iterating this recursion, the error is controlled by the norm of the optimal dual solution. Finally, invoking the uniform bound $\|z^{\star}\|\leq\sqrt{2n}\,R$ leads directly to the quantitative decay rate stated in Lemma 2.
Lemma 2.
Let $\{\check{z}^{t}\}_{t\geq 0}$ be the iterates generated by Algorithm 2, and $z^{\star}$ be an optimal solution of the dual entropic regularized POT problem satisfying $\|z^{\star}\|\leq\sqrt{2n}\,R$. Then we have $\delta_{t}\leq\frac{4nLR^{2}}{(t+1)^{2}}.$
Remark 3.
The $O(1/t^{2})$ rate controls the dual optimality gap, but the algorithm terminates based on the feasibility error $E(t)$. By relating $E(t)$ to the potential decrease in each iteration, one obtains an iteration bound in terms of $L,R,n$ and the target $\varepsilon^{\prime}$, we end up with Theorem 3.
Theorem 3.
Let $\left\{\left(z^{t}\right)\right\}_{t\geq 0}$ be the iterates generated by Algorithm 2. The number of iterations required to reach the stopping criterion $E_{t}\leq\varepsilon^{\prime}$ satisfies $t\leq 1+\left(\frac{12\sqrt{14nL}R}{\varepsilon^{\prime}}\right)^{2/3}.$
Proof sketch.
The proof proceeds by bounding the potential decrease $S=\varphi\left(z^{t}\right)-\varphi\left(\check{z}^{t+1}\right)$. If the update is on $u$ (or symmetrically $v$), one rewrites $S$ in terms of the marginal constraints and shows that it reduces to a sum of divergences $\rho(r_{i},r_{i}(B(z^{t})+\tilde{p}_{i}^{t})$. If the update is on $w$, a similar calculation using the update rule for $w^{t+1}$ shows that $S=\rho(s,\|B(z^{t})\|_{1})$. Combining both cases yields a lower bound for $S$ in terms of all marginal violations, namely $S\geq\tfrac{1}{3}\sum_{i}\rho(\cdot)+\tfrac{1}{3}\sum_{j}\rho(\cdot)+\tfrac{1}{3}\rho(\cdot)$. Next, Lemma 6 of [1], together with Cauchy–Schwarz, shows that $S\geq\tfrac{1}{63}E(t)^{2}$. This gives a telescoping inequality on the dual objective $\varphi$. Using Lemma 2, we obtain $\sum_{i=j}^{t}E_{i}^{2}\leq 252nLR^{2}/(j+1)^{2}$. Since $E_{t}$ must stay above the stopping threshold $\varepsilon^{\prime}$ (line 10 in Algorithm 2), this would lead to a contradiction unless $t$ is bounded. Now choosing $j=t/2$ yields the final rate $t\leq 1+\left(\tfrac{12\sqrt{14nL}R}{\varepsilon^{\prime}}\right)^{2/3}$. ∎
4 Experimental results
4.1 Color transfer
Experimental setup. We follow the setup of [4]. AI-generated images (Fig. 1) are represented in $\mathbb{R}^{3}$ by RGB values of pixels and quantized to $n=800$ colors using $k$ means, producing histograms $a,b$ as source and target distributions. Both are normalized by $\max\{\|a\|_{1},\|b\|_{1}\}$ so total mass $\leq 1$. The cost is $C_{ij}=\|a_{i}-b_{j}\|^{2}/\max(C)$, and the transport budget is $s=0.2\cdot\min(\sum a,\sum b)$. We compare ASPOT, feasible Sinkhorn, and APDAGD, each run for 1500 iterations with $\gamma=10^{-3}$ and $\texttt{tol}=10^{-7}$. After computing a plan $X$, barycentric projection $\hat{a}_{i}=\sum_{j}X_{ij}b_{j}/\sum_{j}X_{ij}$ recolors pixels assigned to centroid $a_{i}$, transforming the new image. ASPOT results. Fig. 1 compares the three methods on color transfer. The convergence curves show that ASPOT rapidly reduces transport cost, reaching good solutions far earlier than the others. Sinkhorn progresses steadily but much slower, while APDAGD stalls for many iterations before improving. As a result, when ASPOT is already near optimal, the other methods still lag by a gap consistent with ASPOT’s better $\varepsilon$-dependence. The visual results confirm that APDAGD and Sinkhorn yield almost identical outputs, both washed out, lacking detail, with incomplete recoloring and visible artifacts, reflecting their slow convergence. By contrast, ASPOT transfers cool tones more clearly and preserves source structure, producing sharper and more natural images. Overall, ASPOT both converges faster and produces better visual quality than the baselines.
Fig. 2: Tuned Sinkhorn Technique Validation
Tuned Sinkhorn results We evaluate tuned Sinkhorn on the color transfer framework described earlier. In this setting, classical Sinkhorn and tuned Sinkhorn technique, use the same tolerance $\texttt{tol}=10^{-2}$, and the scaling parameter A_mult is set to $\|C\|_{\infty}/\texttt{tol}$. For the baseline Sinkhorn, $\gamma$ and stopping rule follow the default choice from (Algorithm 2, Lines 1–8), while for tuned Sinkhorn these parameters are selected according to our theoretical prescription in Theorem 2. Fig. 2 reports $\langle C,X\rangle$ against the number of iterations in log-log scale. The behavior matches our theoretical insights: as $p$ increases, the convergence of tuned Sinkhorn accelerates noticeably, with larger $p$ leading to faster descent. Most importantly, by applying the tuned Sinkhorn technique, it consistently outperforms the classical Sinkhorn across all settings. This empirically validates the sharper complexity bound for fine-tuned Sinkhron in Theorem 2.
4.2 Point Cloud Registration
We follow the protocol of [22, 28], registering two 3-D point clouds by applying translations and rotations to align the source cloud with $m$ points to the target cloud with $n$ points. The transported mass is $s=\alpha\cdot\tfrac{\min\{m,n\}}{\max\{m,n\}}$, with $\alpha$ a tunable constant; smaller values improve convergence. In our setup, the blue cloud initially contains 50% of the red but is misaligned. We set $\alpha=0.4$ for ASPOT and $\alpha=0.7$ for Sinkhorn, with $\gamma=4.4^{-3}$ and annealing rate $0.83$. Different $\alpha$ values are used because Sinkhorn becomes unstable at smaller settings.
Fig. 3: Point Cloud Registration Experiment
Experimental results. ASPOT successfully aligns the two point clouds within 41 registrations, reaching the threshold $10^{-5}$, while Sinkhorn requires 43. Each registration under ASPOT also converges with fewer iterations to compute the POT map. This is illustrated in Fig. 3 where the “Before” and “After” visualizations confirm that the transformed blue cloud overlaps tightly with the red one, demonstrating accurate alignment. The right-hand plot further shows that $\langle C,X\rangle$ under ASPOT decreases in an almost straight line on the log scale, reflecting steady reduction of transport cost, whereas Sinkhorn descends more slowly. Taken together, these results highlight two concrete advantages of ASPOT: fewer registrations overall and faster per-registration convergence.
5 Conclusion
We developed ASPOT and analyze a tuned varriant technique of Sinkhorn through parameter optimization. Theoretically, we derived better complexity bounds by combining Nesterov-style momentum with Greenkhorn updates and by showing how tuning the regularization parameter improves rates of the classical Sinkhorn. Empirically, color transfer and point cloud experiments confirmed these results: ASPOT converges faster than baselines, while the tuned technique for Sinkhorn consistently improves over the classical version, especially as $p$ increases. We hope these findings can help guide future research on scaling POT to larger and multimarginal settings.
6 Acknowledgement
The authors express their profound gratitude to Mr. Quang Minh Nguyen (MIT, nmquang@mit.edu) and Mr. Hoang Huy Nguyen (Georgia Tech, hnguyen455@gatech.edu) for their invaluable input in this project. This research is supported by VNUHCM-University of Information Technology’s Scientific Research Support Fund.